Circular No. 7862

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002bt IN UGC 8584 W. D. Li, University of California at Berkeley, reports the LOTOSS

(cf. IAUC 7514) discovery of an apparent supernova (mag ∼ 17.9) on an unfiltered CCD image taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Mar. 25.4 UT. The new object was confirmed on an image taken in moonlight on Mar. 27.4. SN 2002bt is located at α = 13^{h}36^{m}13^{s}.20, δ = −1^{o}02′19′′.8 (equinox 2000.0), which is 13′′.2 west and 1′′.9 south of the nucleus of the central galaxy of UGC 8584, a triple interacting galaxy system. A KAIT image taken on Jan. 19.5 showed nothing at this position (limiting mag ∼ 19.0).

COMETS C/2002 E3, C/2002 F2, AND C/2002 F3 (SOHO) Further to IAUC 7854, D. Hammer reports his measurements of ad-

ditional Kreutz sungrazing comets found on SOHO website images by X. Leprette (C/2002 E3, C3 coronagraph), S. Hönig (C/2002 F2; C2 and C3), and M. Meyer (C/2002 F3; C2). Reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 E3 Mar. 15.570  0^{h}00.^{m}9  − 2^{o}59′  2002-F35
C/2002 F2  20.321  0 17.5  − 1 16  2002-F43
C/2002 F3  25.204  0 23.4  + 1 22  2002-F61

COMET C/2002 C1 (IKEYA-ZHANG) C. W. Hergenrother, Lunar and Planetary Laboratory, reports that

observations with the Catalina 1.5-m reflector on Mar. 3.11, 4.10, 5.10, 9.10, 10.10, and 26.10 UT by J. Barnes, J. Fortney, and himself, and with a 0.28-m reflector on Mar. 9.81 by G. Masi (Ceccano, Italy), show dust shells within 30′′ of the nucleus of comet C/2002 C1 on the sunward side. The shells display evidence of two discrete jets.

COMET C/2002 F1 (UTSUNOMIYA) Visual m_{1} estimates: Mar. 21.14 UT, 9.5 (M. Reszelski, Szamotuly,

Poland, 0.41-m reflector); 22.20, 9.3 (J. J. Gonzalez, Asturias, Spain, 11×80 binoculars); 23.15, 9.2 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 23.89, 9.5 (A. Pearce, Noble Falls, W. Australia, 0.20-m reflector); 27.21, 8.7 (Gonzalez).

2002 March 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green